# Adaptive Massively Parallel Coloring in Sparse Graphs

### v1

@@ 4.2 Game Analysis
The following lemma was used in Lemma 4.2. Informally, it ensures that more than half of coins forwarded by a node $u\in S_{v}$ are forwarded to nodes in $D(\ell_{\beta},u)$.
@@
For readability, we write $\sigma$ and $\ell$ instead of $\sigma_{S_{v},\beta}$ and $\ell_{\beta}$, respectively. Consider $u\in S_{v}$ as in the lemma statement. Since $\deg(u)\geq\beta+1$, it holds that $|F(\sigma,u)|=\beta+1$ by Definition 4.1. Since $\ell(u)\neq\infty$, it holds that $|N(u)\setminus D(\ell,u)|\leq\beta$ by Lemma 3.11. Recall that $\ell(w)<\ell(u)$ holds for every node $w\in D(\ell,u)\setminus\{u\}$ by Definition 3.9. Hence, it holds that $\ell(w)<\infty$ for every node $w\in D(\ell,u)\setminus\{u\}$. Let us consider the case where $\sigma(u)=\infty$. By Lemma 3.7 there are at least $\beta+1$ nodes $w^{\prime}\in N(u)$ with $\sigma(w^{\prime})=\infty$. Hence, it holds that $\sigma(w)=\infty$ for every node $w\in F(\sigma,u)$. Since $|N(u)\setminus D(\ell,u)|\leq\beta$, $|F(\sigma,u)|=\beta+1$, and $F(\sigma,u)\subseteq N(u)$, there must be at least one node $w\in F(\sigma,u)\cap D(\ell,u)$. Because $w\in F(\sigma,u)$, it holds that $\sigma(w)=\infty$. Because $w\in D(\ell,u)\setminus\{u\}$, it holds that $\ell(w)<\infty$, proving the claim. Let us consider the case where $\sigma(u)\neq\infty$. By Lemma 3.7, it holds that $\sigma(w)\geq\sigma(u)-1$ for at least $\beta+1$ nodes $w\in N(u)$. Hence, it holds that $\sigma(w)\geq\sigma(u)-1$ for every node $w\in F(\sigma,u)$. Since $|N(u)\setminus D(\ell,u)|\leq\beta$, $|F(\sigma,u)|=\beta+1$, and $F(\sigma,u)\subseteq N(u)$, there must be at least one node $w\in F(\sigma,u)\cap D(\ell,u)$. Because $w\in F(\sigma,u)$, it holds that $\sigma(w)\geq\sigma(u)-1\geq\ell(u)$. Because $w\in D(\ell,u)$, it holds that $\ell(u)>\ell(w)$. We conclude that $\sigma(w)>\ell(w)$. ∎
@@
During the game, a node can be added to $S_{v}$ only if it is adjacent to a node already in $S_{v}$. Hence, $G[S_{v}]$ is always connected. Consider one super-iteration of the $(x,\beta,F)$-coin dropping game. Observe that the forwarding set of any node $u\in S_{v}$ is of size $\leq\beta+1$, and only nodes with $\geq\beta+1$ coins forward coins. Hence, if a node $w\not\in S_{v}$ receives coins, it receives $\geq 1$ coins. Since the total number of usable coins in a super-iteration is $x$, the number of nodes that can be added to $S_{v}$ in a super-iteration is at most $x$. The claim follows since there are $x^{2}$ super-iterations and $G[S_{v}]$ is connected. ∎
@@ 4.3 The LCA
For any constant $\varepsilon>0$ and $\beta\geq(2+\varepsilon)\alpha$ there is a deterministic LCA algorithm that uses at most $x^{6}$ queries per node on a graph $G$ with arboricity $\alpha$ and assigns each node a layer from ${}\mathbb{N}\cup\{\infty\}$ such that the following holds: normal-$\triangleright$ There exists a subset $S\subseteq V$ containing at least a $1-2^{1-\log x/\log_{\beta/(2\alpha)}(\beta+1)}$ fraction of vertices such that the layering of $S$ forms a $\beta$-partition of $G[S]$ with $\log_{\beta+1}x$ layers. normal-$\triangleleft$
@@
Consider $\varepsilon$ and $\beta$ as in the lemma statement. When questioned about vertex $v$, the LCA performs the $(x,\beta,F)$-coin dropping game and outputs the $S_{v}$-induced $\beta$-partition layer of $v$. In order to perform the coin dropping game, the LCA needs to be able to compute the $S_{v}$-induced $\beta$-partition (Definition 3.6) locally. This is possible since it only requires knowing $G[S_{v}]$ and the degrees of nodes in $S_{v}$. By Lemma 4.6, subgraph $G[S_{v}]$ contains at most $x^{6}$ edges at any point during the game, bounding the query complexity of the LCA. By Lemma 4.4, if $|D(\ell_{\beta},v)|\leq x^{2}$ and $\ell_{\beta}(v)\leq\log_{\beta+1}x$, then upon termination the LCA can output $\ell_{\beta}(v)$, which is the layer of $v$ in the natural $\beta$-partition of graph $G$. By Lemma 4.9, there are at least $|V|\cdot(1-2^{1-\log x/\log_{\beta/(2\alpha)}(\beta+1)})$ such nodes, completing the proof. ∎
@@
Instead of Lemma 4.7, we actually obtain a stronger statement. Our LCA, if queried about a node $u$, outputs a partial $\beta$-partition $\ell_{u}:V\rightarrow\mathbb{N}\cup\{\infty\}$. normal-$\triangleright$ There exists a subset $S\subseteq V$ containing at least a $1-2^{1-\log x/\log_{\beta/(2\alpha)}(\beta+1)}$ fraction of vertices, and $\lambda:V\rightarrow\mathbb{N}\cup\{\infty\}:v\mapsto\min_{u\in S}\ell_{u}(v)$ is a partial $\beta$-partition of $G$ with $\log_{\beta+1}x$ layers normal-$\triangleleft$
@@
In order to prove that $\lambda$ is a partial $\beta$-partition, we need to show that $\lambda(w)\geq\lambda(v)\neq\infty$ for at most $\beta$ nodes $w\in N(v)$. Let $N^{\geq}(v)$ be the set of neighbors $w\in N(v)$ with $\lambda(w)\geq\lambda(v)$. Without loss of generality, we can assume that $\lambda(v)=\ell_{1}(v)$. Let $N^{<}_{1}(v)$ be the set of neighbors $w\in N(v)$ with $\ell_{1}(w)<\ell_{1}(v)$, and let $N^{\geq}_{1}(v)$ be the set of neighbors $w\in N(v)$ with $\ell_{1}(w)\geq\ell_{1}(v)$. Let us count how many nodes from sets $N^{<}_{1}(v)$ and $N^{\geq}_{1}(v)$ could be in $N^{\geq}(v)$.
@@ 5 $\beta$-partition in AMPC
Consider an graph $G=(V,E)$ with $n$ nodes and arboricity $\alpha<n^{\frac{(\delta/c)^{2}\cdot\log((2+\varepsilon)/2)}{2(2+\varepsilon)}}$ (this requirement is lifted later in the proof). The algorithm operates on data stores $D_{0},D_{1},\dots$ such that the total number of data stores equals the number of rounds of the algorithm. Let us keep track of two types of key-value pairs in every data store $D_{i}$. The first type being a value from ${}\mathbb{N}\cup\{\infty\}$ returned for every node $v\in V$, which represents $v$’s layer in a (partial) $\beta$-partition. The second type is defined using the first type as follows. Let $G_{i}=(V_{i},E_{i})$ be the subgraph of $G=(V,E)$, induced by nodes $v$ with $D_{i}(v)=\infty$, which has arboricity at most $\alpha$ by Fact 3.2. The edges of graph $G_{i}$ are stored in the data store $D_{i}$ as key-value pairs $(v,\deg_{G_{i}}(v),j)\xrightarrow{}(u,\deg_{G_{i}}(u))$, where $j\in[0,\deg_{G_{i}}(v)-1]$ and node $u$ is the $j$th neighbor of $v$ in $G_{i}$, according to some arbitrary ordering. For $D_{0}$, the values of the first type are set to $\infty$, and hence $G_{0}=G$ is just the input graph. Let $P=n$ be the number of ${}^{+}$19] provides the necessary details for employing this number of machines using parallel slackness. and $S=\Omega(n^{\delta})$ be the local space of a machine, for any $\delta>0$. In order to compute $D_{i+1}$ from $D_{i}$, we have the following protocol. Let $M_{v}$ be the machine responsible for node $v\in V$. Using $D_{i}$, or more precisely graph $G_{i}$, machine $M_{v}$ performs the LCA of Remark 4.8. The query complexity of the LCA is $x^{6}$, so when $x=n^{\delta/c}$ for some $c>6$, the local and total space of our AMPC algorithm are respected. Next, we need to argue that the partial $\beta$-partition of $G_{i}$ from Remark 4.8 is computable. After performing the LCA from Remark 4.8, every machine $M_{v}$ has the $S_{v}$-induced $\beta$-partition $\sigma_{S_{v},\beta}$ stored in memory. Machine $M_{v}$ can write the key-value pair $u\xrightarrow{}\sigma_{S_{v},\beta}(u)$ for every node $u\in S_{v}$ for which $\sigma_{S_{v},\beta}(u)\leq\log_{\beta+1}x$ to data store $D_{i+1}$. The separate set of machines that handles the DDS sorts through all written values and keeps the smallest $\sigma(u)$ value for every node $u\in V_{i}$. Observe that in order to make the (final outputted) $\beta$-partition consistent throughout the algorithm, we need to record $\sigma(u)$ plus the maximum numbered layer of any node in $D_{i}$, effectively appending the layers. Additionally, using $D_{i}$, these machines can compute $\deg_{G_{i+1}}(u)$ for every node $u\in G_{i+1}$ and port the edges of $G_{i+1}$ to $D_{i+1}$. Let us analyze the size of $|V_{i}|$. By applying Remark 4.8 for every rounds, it holds that
@@
*$\stackrel{{\scriptstyle*}}{{\leq}}\frac{\log(2(2+\varepsilon)\alpha)}{\log((2+\varepsilon)\alpha)-\log 2\alpha}-\delta/c\cdot\log n$
@@
where in *** we use $\beta\geq(2+\varepsilon)\alpha$. Let $d\coloneqq(\delta/c-(\delta/c)^{2})$. Observe that since $\delta/c<1$, it holds that $d>0$. Plugging Equation 2 into Equation 1 gives
@@
By Remark 4.8, the maximum layer in the partial $\beta$-partition computed for every data store is $\log_{\beta+1}x$. Hence, after $O(\log_{\beta/(2\alpha)}\beta)$ rounds, the size of the $\beta$-partition is $=O(\log_{\beta/(2\alpha)}n)$. When $\alpha\geq n^{\frac{(\delta/c)^{2}\cdot\log((2+\varepsilon)/2)}{2(2+\varepsilon)}}$, we employ a different algorithm altogether. We simulate the $H$-partition algorithm of [BE10b]. For every data store $D_{i}$, instead of performing the $(x,\beta,F)$-coin dropping game, every node of degree $\leq\beta$ simply places itself in layer $i$. By Lemma 3.4, for any $\beta\geq(2+\varepsilon)\alpha$, this algorithm results in a $\beta$-partition of size $O(\log_{\beta/(2\alpha)}n)$ in $O(\log_{\beta/(2\alpha)}n)=O(\log_{\beta/(2\alpha)}\beta)$ time. ∎
@@
Let $\beta$ be defined using arboricity $\alpha$. When $\alpha$ is unknown, we can still find the $\beta$-partition with no additional cost in the runtime or memory by employing a guessing scheme for $\alpha$. First, we perform a series of sequential executions of Theorem 1.2, such that instance $i$ computes a $\beta_{i}$-partition, using the $i$th guess for arboricity $\alpha_{i}=2^{2^{i}}$, $i\in[0,\log\log n]$. We terminate after the first guess $\alpha_{k}$ which returns a $\beta_{k}$-partition. Due to the double exponential guessing and the fact that a correct guess we yield partition, we know that $a_{k}<\alpha^{2}$. Next, we perform a number of parallel executions of Theorem 1.2, such that instance $i$ computes a $\beta_{i}$-partition, where the $i$th guess for arboricity is $\alpha_{i}=\sqrt{a_{k}}\cdot(1+\varepsilon)^{i}$, $i\in[0,\log_{1+\varepsilon}\sqrt{a_{k}}]$ for some $\varepsilon>0$. Observe that for some instance $j$, it must hold that $\alpha_{j}<(1+\varepsilon)\alpha$, producing a $\beta_{i}$-partition of size $O(\log_{\beta_{i}/2\alpha_{i}}n)=O(\log_{\beta/2\alpha}n)$. As our solution, we pick the smallest instance that produces a partition. Let us analyze the total runtime when employing the guessing scheme above. The sequential executions take
@@
rounds in total. The parallel executions take $O(\log_{\beta/2\alpha}\alpha^{2})=O(\log_{\beta/2\alpha}\beta)$ rounds. The parallel executions require an additional $\log_{1+\varepsilon}\sqrt{a_{k}}=O(\log n)$ factor of space, which can be allowed by scaling the constant $\delta$, corresponding to the local space $O(n^{\delta})$. It is worth mentioning that since $\alpha_{j}$ may differ from $\alpha$ by a $(1+\varepsilon)$ factor, the resulting $\beta_{j}$ may also differ from $\beta$ by some $\varepsilon^{\prime}>1$ factor. This can be mended by simply using a slightly smaller $\beta$ value to begin with, which is always possible due to the slack in requirement $\beta\geq(2+\epsilon)\alpha,\leavevmode\nobreak\ \epsilon>0$. ∎

### v2

@@ 4.2 Game Analysis
The following lemma was used in Lemma 4.2. Informally, it ensures that roughly a $1/\beta$ fraction of coins forwarded by a node $u\in S_{v}$ are forwarded to nodes in $D(\ell_{\beta},u)$.
@@
For readability, we write $\sigma$ and $\ell$ instead of $\sigma_{S_{v},\beta}$ and $\ell_{\beta}$, respectively. Consider $u\in S_{v}$ as in the lemma statement. Since $\deg(u)\geq\beta+1$, it holds that $|F(\sigma,u)|=\beta+1$ by Definition 4.1. Since $\ell(u)\neq\infty$, it holds that $|N(u)\setminus D(\ell,u)|\leq\beta$ by Lemma 3.11. Recall that $\ell(w)<\ell(u)$ holds for every node $w\in D(\ell,u)\setminus\{u\}$ by Definition 3.9. Hence, it holds that $\ell(w)<\infty$ for every node $w\in D(\ell,u)\setminus\{u\}$.
Let us consider the case where $\sigma(u)=\infty$. By Lemma 3.7 there are at least $\beta+1$ nodes $w^{\prime}\in N(u)$ with $\sigma(w^{\prime})=\infty$. Hence, it holds that $\sigma(w)=\infty$ for every node $w\in F(\sigma,u)$. Since $|N(u)\setminus D(\ell,u)|\leq\beta$, $|F(\sigma,u)|=\beta+1$, and $F(\sigma,u)\subseteq N(u)$, there must be at least one node $w\in F(\sigma,u)\cap D(\ell,u)$. Because $w\in F(\sigma,u)$, it holds that $\sigma(w)=\infty$. Because $w\in D(\ell,u)\setminus\{u\}$, it holds that $\ell(w)<\infty$, proving the claim.
Let us consider the case where $\sigma(u)\neq\infty$. By Lemma 3.7, it holds that $\sigma(w)\geq\sigma(u)-1$ for at least $\beta+1$ nodes $w\in N(u)$. Hence, it holds that $\sigma(w)\geq\sigma(u)-1$ for every node $w\in F(\sigma,u)$. Since $|N(u)\setminus D(\ell,u)|\leq\beta$, $|F(\sigma,u)|=\beta+1$, and $F(\sigma,u)\subseteq N(u)$, there must be at least one node $w\in F(\sigma,u)\cap D(\ell,u)$. Because $w\in F(\sigma,u)$, it holds that $\sigma(w)\geq\sigma(u)-1\geq\ell(u)$. Because $w\in D(\ell,u)$, it holds that $\ell(u)>\ell(w)$. We conclude that $\sigma(w)>\ell(w)$. ∎
@@
During the game, a node can be added to $S_{v}$ only if it is adjacent to a node already in $S_{v}$. Hence, $G[S_{v}]$ is always connected.
Consider one super-iteration of the $(x,\beta,F)$-coin dropping game. Observe that the forwarding set of any node $u\in S_{v}$ is of size $\leq\beta+1$, and only nodes with $\geq\beta+1$ coins forward coins. Hence, if a node $w\not\in S_{v}$ receives coins, it receives $\geq 1$ coins. Since the total number of usable coins in a super-iteration is $x$, the number of nodes that can be added to $S_{v}$ in a super-iteration is at most $x$. The claim follows since there are $x^{2}$ super-iterations and $G[S_{v}]$ is connected. ∎
@@ 4.3 The LCA
For any constant $\varepsilon>0$ and $\beta\geq(2+\varepsilon)\alpha$ there is a deterministic LCA algorithm that uses at most $x^{6}$ queries per node on a graph $G$ with arboricity $\alpha$ and assigns each node a layer from ${}\mathbb{N}\cup\{\infty\}$ such that the following holds:
$\triangleright$ There exists a subset $S\subseteq V$ containing at least a $1-2^{1-\log x/\log_{\beta/(2\alpha)}(\beta+1)}$ fraction of vertices such that the layering of $S$ forms a $\beta$-partition of $G[S]$ with $\log_{\beta+1}x$ layers. $\triangleleft$
@@
Consider $\varepsilon$ and $\beta$ as in the lemma statement. When questioned about vertex $v$, the LCA performs the $(x,\beta,F)$-coin dropping game and outputs the $S_{v}$-induced $\beta$-partition layer of $v$. In order to perform the coin dropping game, the LCA needs to be able to compute the $S_{v}$-induced $\beta$-partition (Definition 3.6) locally. This is possible since it only requires knowing $G[S_{v}]$ and the degrees of nodes in $S_{v}$. By Lemma 4.6, subgraph $G[S_{v}]$ contains at most $x^{6}$ edges at any point during the game, bounding the query complexity of the LCA.
By Lemma 4.4, if $|D(\ell_{\beta},v)|\leq x^{2}$ and $\ell_{\beta}(v)\leq\log_{\beta+1}x$, then upon termination the LCA can output $\ell_{\beta}(v)$, which is the layer of $v$ in the natural $\beta$-partition of graph $G$. By Lemma 4.9, there are at least $|V|\cdot(1-2^{1-\log x/\log_{\beta/(2\alpha)}(\beta+1)})$ such nodes, completing the proof. ∎
@@
Instead of Lemma 4.7, we actually obtain a stronger statement. Our LCA, if queried about a node $u$, outputs a partial $\beta$-partition $\ell_{u}:V\rightarrow\mathbb{N}\cup\{\infty\}$.
$\triangleright$ There exists a subset $S\subseteq V$ containing at least a $1-2^{1-\log x/\log_{\beta/(2\alpha)}(\beta+1)}$ fraction of vertices, and $\lambda:V\rightarrow\mathbb{N}\cup\{\infty\}:v\mapsto\min_{u\in S}\ell_{u}(v)$ is a partial $\beta$-partition of $G$ with $\log_{\beta+1}x$ layers. $\triangleleft$
@@
In order to prove that $\lambda$ is a partial $\beta$-partition, we need to show that $\lambda(w)\geq\lambda(v)\neq\infty$ for at most $\beta$ nodes $w\in N(v)$. Let $N^{\geq}(v)$ be the set of neighbors $w\in N(v)$ with $\lambda(w)\geq\lambda(v)$.
Without loss of generality, we can assume that $\lambda(v)=\ell_{1}(v)$. Let $N^{<}_{1}(v)$ be the set of neighbors $w\in N(v)$ with $\ell_{1}(w)<\ell_{1}(v)$, and let $N^{\geq}_{1}(v)$ be the set of neighbors $w\in N(v)$ with $\ell_{1}(w)\geq\ell_{1}(v)$. Let us count how many nodes from sets $N^{<}_{1}(v)$ and $N^{\geq}_{1}(v)$ could be in $N^{\geq}(v)$.
@@ 5 $\beta$-partition in AMPC
Consider an graph $G=(V,E)$ with $n$ nodes and arboricity $\alpha<n^{\frac{(\delta/c)^{2}\cdot\log((2+\varepsilon)/2)}{2(2+\varepsilon)}}$ (this requirement is lifted later in the proof). The algorithm operates on data stores $D_{0},D_{1},\dots$ such that the total number of data stores equals the number of rounds of the algorithm. Let us keep track of two types of key-value pairs in every data store $D_{i}$. The first type being a value from ${}\mathbb{N}\cup\{\infty\}$ returned for every node $v\in V$, which represents $v$’s layer in a (partial) $\beta$-partition. The second type is defined using the first type as follows. Let $G_{i}=(V_{i},E_{i})$ be the subgraph of $G=(V,E)$, induced by nodes $v$ with $D_{i}(v)=\infty$, which has arboricity at most $\alpha$ by Fact 3.2. The edges of graph $G_{i}$ are stored in the data store $D_{i}$ as key-value pairs $(v,\deg_{G_{i}}(v),j)\xrightarrow{}(u,\deg_{G_{i}}(u))$, where $j\in[0,\deg_{G_{i}}(v)-1]$ and node $u$ is the $j$th neighbor of $v$ in $G_{i}$, according to some arbitrary ordering. For $D_{0}$, the values of the first type are set to $\infty$, and hence $G_{0}=G$ is just the input graph.
Let $P=n$ be the number of machines444[BDE+19] provides the necessary details for employing this number of machines using parallel slackness. and $S=\Omega(n^{\delta})$ be the local space of a machine, for any $\delta>0$. In order to compute $D_{i+1}$ from $D_{i}$, we have the following protocol. Let $M_{v}$ be the machine responsible for node $v\in V$. Using $D_{i}$, or more precisely graph $G_{i}$, machine $M_{v}$ performs the LCA of Remark 4.8. The query complexity of the LCA is $x^{6}$, so when $x=n^{\delta/c}$ for some $c>6$, the local and total space of our AMPC algorithm are respected.
Next, we need to argue that the partial $\beta$-partition of $G_{i}$ from Remark 4.8 is computable. After performing the LCA from Remark 4.8, every machine $M_{v}$ has the $S_{v}$-induced $\beta$-partition $\sigma_{S_{v},\beta}$ stored in memory. Machine $M_{v}$ can write the key-value pair $u\xrightarrow{}\sigma_{S_{v},\beta}(u)$ for every node $u\in S_{v}$ for which $\sigma_{S_{v},\beta}(u)\leq\log_{\beta+1}x$ to data store $D_{i+1}$. The separate set of machines that handles the DDS sorts through all written values and keeps the smallest $\sigma(u)$ value for every node $u\in V_{i}$. Observe that in order to make the (final outputted) $\beta$-partition consistent throughout the algorithm, we need to record $\sigma(u)$ plus the maximum numbered layer of any node in $D_{i}$, effectively appending the layers. Additionally, using $D_{i}$, these machines can compute $\deg_{G_{i+1}}(u)$ for every node $u\in G_{i+1}$ and port the edges of $G_{i+1}$ to $D_{i+1}$.
Let us analyze the size of $|V_{i}|$. By applying Remark 4.8 for every rounds, it holds that
@@
$\stackrel{{\scriptstyle*}}{{\leq}}\frac{\log(2(2+\varepsilon)\alpha)}{\log((2+\varepsilon)\alpha)-\log 2\alpha}-\delta/c\cdot\log n$
@@
where in * we use $\beta\geq(2+\varepsilon)\alpha$. Let $d\coloneqq(\delta/c-(\delta/c)^{2})$. Observe that since $\delta/c<1$, it holds that $d>0$. Plugging Equation 2 into Equation 1 gives
@@
By Remark 4.8, the maximum layer in the partial $\beta$-partition computed for every data store is $\log_{\beta+1}x$. Hence, after $O(\log_{\beta/(2\alpha)}\beta)$ rounds, the size of the $\beta$-partition is $=O(\log_{\beta/(2\alpha)}n)$.
When $\alpha\geq n^{\frac{(\delta/c)^{2}\cdot\log((2+\varepsilon)/2)}{2(2+\varepsilon)}}$, we employ a different algorithm altogether. We simulate the $H$-partition algorithm of [BE10b]. For every data store $D_{i}$, instead of performing the $(x,\beta,F)$-coin dropping game, every node of degree $\leq\beta$ simply places itself in layer $i$. By Lemma 3.4, for any $\beta\geq(2+\varepsilon)\alpha$, this algorithm results in a $\beta$-partition of size $O(\log_{\beta/(2\alpha)}n)$ in $O(\log_{\beta/(2\alpha)}n)=O(\log_{\beta/(2\alpha)}\beta)$ time. ∎
@@
Let $\beta$ be defined using arboricity $\alpha$. When $\alpha$ is unknown, we can still find the $\beta$-partition with no additional cost in the runtime or memory by employing a guessing scheme for $\alpha$.
First, we perform a series of sequential executions of Theorem 1.2, such that instance $i$ computes a $\beta_{i}$-partition, using the $i$th guess for arboricity $\alpha_{i}=2^{2^{i}}$, $i\in[0,\log\log n]$. We terminate after the first guess $\alpha_{k}$ which returns a $\beta_{k}$-partition. Due to the double exponential guessing and the fact that a correct guess we yield partition, we know that $a_{k}<\alpha^{2}$.
Next, we perform a number of parallel executions of Theorem 1.2, such that instance $i$ computes a $\beta_{i}$-partition, where the $i$th guess for arboricity is $\alpha_{i}=\sqrt{a_{k}}\cdot(1+\varepsilon)^{i}$, $i\in[0,\log_{1+\varepsilon}\sqrt{a_{k}}]$ for some $\varepsilon>0$. Observe that for some instance $j$, it must hold that $\alpha_{j}<(1+\varepsilon)\alpha$, producing a $\beta_{i}$-partition of size $O(\log_{\beta_{i}/2\alpha_{i}}n)=O(\log_{\beta/2\alpha}n)$. As our solution, we pick the smallest instance that produces a partition.
Let us analyze the total runtime when employing the guessing scheme above. The sequential executions take
@@
rounds in total. The parallel executions take $O(\log_{\beta/2\alpha}\alpha^{2})=O(\log_{\beta/2\alpha}\beta)$ rounds. The parallel executions require an additional $\log_{1+\varepsilon}\sqrt{a_{k}}=O(\log n)$ factor of space, which can be allowed by scaling the constant $\delta$, corresponding to the local space $O(n^{\delta})$.
It is worth mentioning that since $\alpha_{j}$ may differ from $\alpha$ by a $(1+\varepsilon)$ factor, the resulting $\beta_{j}$ may also differ from $\beta$ by some $\varepsilon^{\prime}>1$ factor. This can be mended by simply using a slightly smaller $\beta$ value to begin with, which is always possible due to the slack in requirement $\beta\geq(2+\epsilon)\alpha,\leavevmode\nobreak\ \epsilon>0$. ∎